# LocalSearchBench: Benchmarking Agentic Search in Real-World Local Life Services

### v1

@@ Abstract.
Recent advances in large reasoning models (LRMs) have enabled agentic search systems to perform complex multi-step reasoning across multiple sources. However, most studies focus on general information retrieval and rarely explores vertical domains with unique challenges. In this work, we focus on local life services and introduce LocalSearchBench, which encompass diverse and complex business scenarios. Real-world queries in this domain are often ambiguous and require multi-hop reasoning across merchants and products, remaining challenging and not fully addressed. As the first comprehensive benchmark for agentic search in local life services, LocalSearchBench includes over 150,000 high-quality entries from various cities and business types. We construct 300 multi-hop QA tasks based on real user queries, challenging agents to understand questions and retrieve information in multiple steps. We also developed LocalPlayground, a unified environment integrating multiple tools for agent interaction. Experiments show that even state-of-the-art LRMs struggle on LocalSearchBench: the best model (DeepSeek-V3.1) achieves only 34.34% correctness, and most models have issues with completeness (average 77.33%) and faithfulness (average 61.99%). This highlights the need for specialized benchmarks and domain-specific agent training in local life services. Code, Benchmark, and Leaderboard are available at localsearchbench.github.io
@@ 1. Introduction
Recent advances in large reasoning models (LRMs) have significantly enhanced the capabilities of AI agents, enabling them to perform complex reasoning and planning tasks (xu2025towards; ferrag2025llm; wu2025agentic; li2025webthinker; ke2025survey; chai2025rlfactory). This progress has driven the development of agentic search systems, which can decompose complex queries, execute multi-step reasoning, and dynamically orchestrate multiple information sources to achieve autonomous cross-domain reasoning (sun2025simpledeepsearch; zhang2025evolvesearch; shi2025pangudeepdiveradaptivesearch; shi2025toollearningwildempowering; zhang2025web; li2025search; mo2025conversational; yue2025promoting; yue2025uiorchestra). These systems also demonstrate the potential to expand from general domains to vertical application scenarios such as e-commerce (zhou2024webarenarealistic; yao2023webshopscalable). However, there has been little work applying agentic search to the local life service domain, and the lack of suitable evaluation benchmarks further constrains new research in this area.
@@
Local life services cover a wide range of sub-domains, including dining, lifestyle, shopping, accommodation, travel, and healthcare (lan2025benchmarking; 10.1145/3580305.3599874). Leveraging location-based digital platforms, these services connect users with nearby merchants and service providers, forming complex networks of businesses, user preferences, and location-dependent service offerings, and creating a vast market (10.1145/3534678.3539122; kusk2025flexible; liu2025mrgrp). As shown in Figure 1(1(c)), real-world user queries often involve multi-constraint merchant recommendation, spatiotemporal service chain planning, and event-driven service bundling, which require complex multi-step processes to deliver the final response. lan2025benchmarking target at this domain and propose the first systematic framework, but is limited to single-hop queries and cannot fully reflect the challenges of real business scenarios. Therefore, there is an urgent need for a multi-hop benchmark tailored to local life scenarios, accurately reflecting the complexity of real-world search tasks and fully challenging existing agentic search methods.
In this paper, we propose an agentic search framework LocalPlayground to evaluate LRMs for local life service. Given the absence of appropriate benchmarks and the critical need for comprehensive evaluation systems, our approach addresses these challenges through comprehensive dataset construction, systematic evaluation framework and targeted model evaluation.
We introduce LocalSearchBench, a comprehensive benchmark comprising over 150,000 data entries across 6 categories: dining, shopping, accommodation, travel, and healthcare, spanning 3 major cities in China. These entries are transformed from raw data into high-quality augmented and anonymized datasets.
@@
Based on these data, we propose a high-quality instruction synthesis method to design 300 multi-hop QA tasks across three reasoning dimensions: multi-constraint merchant recommendation, spatiotemporal service chain planning, and event-driven service bundling. These multi-hop QA tasks are constructed through a multi-agent data generation process, requiring agents to perform complex reasoning across multiple information sources and reasoning steps.
@@
We build a comprehensive evaluation environment with RAG based on LocalSearchBench data and web search tools to simulate real-world local life service search scenarios where models need to retrieve and synthesize information from multiple sources, and conduct evaluations of multiple LRMs.
@@ 2.1. Agentic Search
gou2025mind2web2evaluatingagentic define agentic search as systems capable of integrating multiple auxiliary tools to autonomously perform interactive generation through search tools (e.g., web search APIs or external databases) for handling complex search tasks. Leveraging the powerful language understanding capabilities of LLMs, these search agents can proactively address ambiguous or underspecified queries through clarification or reasonable assumptions (yang2025agentic). They decompose the initial search task into manageable sub-queries, perform dynamic reasoning and planning based on accumulated contextual information, adjust search strategies in real time, and synthesize information from multiple sources (li2025towards).
Recent advances in the reasoning abilities of LRMs further foster the development of LRM-based agentic search systems. The most representative products include OpenAI Deep Research (jaech2024openai), Gemini Deep Research (comanici2025gemini), and Tongyi Deep Research (tongyidr). In addition to accessing real-time information through search APIs, these systems leverage various auxiliary tools to enable deeper and broader research on highly complex problems. In the open-source community,  li2025search first proposed integrating the agentic search workflow into the o1-style reasoning process of LRMs to achieve autonomous knowledge supplementation.  li2025webthinker,  jin2025search, and  song2025r1 further enhanced the autonomous search capabilities of LLMs by applying reinforcement learning, enabling the models to generate queries and invoke external retrieval systems during reasoning.
@@ 2.2. Benchmarking Agentic capabilities
Existing benchmarks for agentic search systems in general domain can be broadly categorized into closed-ended and open-ended QA tasks. Closed-ended QA benchmarks include multi-hop reasoning datasets that require synthesizing information from multiple sources (yang2018hotpotqa; ho2020constructing; press2022measuring; trivedi2022musique), challenging QA benchmarks designed for long-horizon questions with long-tail knowledge (wei2025browsecomps; zhou2025browsecompzhbenchmarking; xi2025infodeepseek; pham2025sealqa), and fact-checking tasks that evaluate claim verification abilities (wei2024long; wadden2020fact; jiang2020hover; wang2024mfc). Open-ended QA benchmarks focus on deep information seeking tasks, targeting multi-perspective queries (rosset2024researchy) and expert-level research tasks (du2025deepresearch; bosse2025deep; tan2024proxyqa), with some incorporating multi-modal capabilities (yuan2025videodeepresearch).
@@ 3. Construction of LocalSearchBench
Figure 2. Workflow of Local Merchant Database Construction
@@ 3.1. Local Merchant Database Construction
In this subsection, we present the construction of the local merchant database for LocalSearchBench, which serves as the foundation for multi-hop QA tasks. As shown in Figure 2, our database construction consists of two stages: Merchant Seed Data Collection and Data Augmentation and Synthesis.
@@ 3.1.1. Merchant Seed Data Collection
To ensure LocalSearchBench reflects real-world local life service complexities, we built our dataset based on source data from a leading local life service platform M. We sampled over 260,000 real merchant records spanning from January 2025 to December 2025, covering diverse service categories and geographical distributions across three major Chinese cities, with the generation of these seed data requiring 114 hours of computation time.
@@
Based on platform M’s core business verticals, we identified six primary service categories reflecting real-world user demand: Dining (35%), Lifestyle (25%), Shopping (20%), Accommodation (10%), Healthcare (5%), and Tourism (5%), as shown in Figure 11. We curated a keyword pool of 411 unique terms across these categories to reflect authentic user search behaviors and regional preferences.
@@
The collection spans 3 major Chinese cities with comprehensive coverage of benchmarks (Table 1). The location database includes 859 landmarks across diverse categories: commercial centers, transportation hubs, cultural landmarks, educational institutions, medical facilities, and residential areas. Figure 3 shows word clouds of landmark distributions across the 3 cities.
Table 1. Geographic coverage across 3 major Chinese cities
City Landmarks Districts
Shanghai 295 16
Beijing 287 16
Guangzhou 277 11
Total 859 43
@@
Merchant Seed Data Retrieval Query Generation
The system employed scenario-specific keyword pools, constructing queries through random combinations of 1-2 keywords to simulate natural search patterns. This approach generated diverse query variations from the 411-keyword foundation, ensuring comprehensive coverage of local search intents across regional preferences.
This multi-city keyword strategy captured regional variations across 43 administrative divisions, providing robust coverage of China’s diverse local search ecosystems as the foundation for LocalSearchBench database synthesis.
@@ 3.1.2. Data Augmentation and Synthesis
The original merchant seed data contains only 12 basic fields with many missing or unstandardized attributes, which is inadequate for agentic search tasks that require dense, high-coverage merchant profiles. We employ a Data Augmentation Agent to expand these 12-field records into 29-field profiles using LLMs across six dimensions (see Figure 2 and Figure 12). The agent enriches each merchant’s basic, locational, operational, service, contact, and promotional information through 17 newly added fields, as illustrated in Figure 2. The enhanced dataset covers: (1) basic information with standardized categories and descriptions; (2) location intelligence including structured geographical data, coordinates, and landmarks; (3) business operations such as operating hours and pricing; (4) service characteristics like facility lists; (5) contact information, including phone numbers and emails; and (6) promotional content, such as product descriptions and discount offers.
@@
To comply with data protection regulations and enable safe public release, we employ a Privacy Rewriting Agent on the augmented data (see Figure 13). Real-world merchant data contains personally identifiable information and business-sensitive details that must be anonymized. The agent rewrites 6 privacy-related fields to anonymize merchant identity, location, and contact details while keeping the remaining business information usable, as illustrated in Figure 2.
@@ 3.1.3. Data Validation
It evaluates augmented Data quality against original seed data across six core dimensions: (1) Data Completeness (20% weight) assessing field coverage and information richness; (2) Logical Consistency (30% weight) validating business logic coherence and field correlations including name format compliance, category alignment, and pricing rationality; (3) Geographic Accuracy (10% weight) verifying location information correctness and address format consistency; (4) Content Quality (20% weight) evaluating textual descriptions, language expression, and business relevance; (5) Format Standardization (10% weight) ensuring data format compliance and structural consistency; and (6) Innovation Degree (10% weight) measuring content creativity and merchant name diversity through batch-level uniqueness analysis. Each evaluation metric corresponds to specific field comparisons as detailed in Table 2(4).
Privacy Rewriting Data Validation
It evaluates privacy rewriting data quality against augmented data across four core dimensions: (1) Anonymization Degree (50% weight) ensuring complete merchant identity anonymization including name deidentification, location sanitization, and contact information anonymization; (2) Logical Consistency (20% weight) maintaining field relationships and business logic coherence after privacy rewriting; (3) Content Consistency (20% weight) preserving non-sensitive information alignment with previous stage outputs; and (4) Innovation Degree (10% weight) assessing the creativity and diversity of anonymized merchant names to prevent pattern recognition. Each evaluation metric corresponds to specific field comparisons as detailed in Table 2(4).
@@ 3.1.4. Local Merchant Database Statistics
Overall, we curated 150,031 representative samples from a pool of over 260,000 original seed records to create synthesized data, which constitute our LocalSearchBench dataset.
The proportions of each category, the number of data for each city are presented in Figure 4. The distribution closely mirrors that of the source data, ensuring our benchmark maintains the same categorical balance as real-world local life service platforms. The geographical distribution of merchant data for each city was derived through geospatial inference to generate latitude-longitude heatmaps, as shown in Figure 5. The distribution patterns align well with urban clustering phenomena and regional economic hierarchies, demonstrating the dataset’s realistic geographical representation.
(a) City distribution
(b) Shanghai
(c) Beijing
(d) Guangzhou
Figure 4. City distribution and category distribution of 150,031 merchant data across three cities.
@@
Figure 5. Geographical distribution heatmaps of 150,031 merchant data across three cities.

### v2

@@ Abstract.
Recent advances in large reasoning models (LRMs) have enabled agentic search systems to perform complex multi-step reasoning across multiple sources. However, most studies focus on general information retrieval and rarely explores vertical domains with unique challenges. In this work, we focus on local life services and introduce LocalSearchBench, which encompass diverse and complex business scenarios. Real-world queries in this domain are often ambiguous and require multi-hop reasoning across merchants and products, remaining challenging and not fully addressed. As the first comprehensive benchmark for agentic search in local life services, LocalSearchBench comprises a database of over 1.3M merchant entries across 6 service categories and 9 major cities, and 900 multi-hop QA tasks from real user queries that require multi-step reasoning. We also developed LocalPlayground, a unified environment integrating multiple tools for LRMs interaction. Experiments show that even state-of-the-art LRMs struggle on LocalSearchBench: the best model (DeepSeek-V3.2) achieves only 35.60% correctness, and most models have issues with completeness (average 60.32%) and faithfulness (average 30.72%). This highlights the need for specialized benchmarks and domain-specific agent training in local life services. Code, Benchmark, and Leaderboard are available at localsearchbench.github.io.
@@ 1. Introduction
Recent advances in large reasoning models (LRMs) have significantly enhanced the capabilities of AI agents, enabling them to perform complex reasoning and planning tasks (Xu et al., 2025; Ferrag et al., 2025; Wu et al., 2025; Li et al., 2025b; Ke et al., 2025; Chai et al., 2025; Chen et al., 2025). This progress has driven the development of agentic search systems, which can decompose complex queries, execute multi-step reasoning, and dynamically orchestrate multiple information sources to achieve autonomous cross-domain reasoning (Sun et al., 2025; Zhang et al., 2025a; Shi et al., 2025a, b; Zhang et al., 2025b; Li et al., 2025a; Mo et al., 2025; Yue et al., 2025b, a). These systems also demonstrate the potential to expand from general domains to vertical application scenarios such as e-commerce (Zhou et al., 2024; Yao et al., 2023). However, there has been little work applying agentic search to the local life service domain, and the lack of suitable evaluation benchmarks further constrains new research in this area.
@@
Local life services are online-offline integrated services connecting local merchants with nearby users, including dining, lifestyle, shopping, accommodation, travel, healthcare and more to meet residents’ daily needs (Lan et al., 2025b, 2023). Leveraging location-based digital platforms, these services form complex networks of businesses, user preferences, and location-dependent service offerings, creating a vast market (Li et al., 2022; Kusk, 2025; Liu et al., 2025a). As shown in Figure 1(1(c)), real-world user queries often involve multi-constraint merchant recommendation, spatiotemporal service chain planning, and event-driven service bundling, which require complex multi-step processes to deliver the final response.
Lan et al. (2025b) target at this domain and propose the first systematic framework, but is limited to single-hop queries and cannot fully reflect the challenges of real business scenarios. Moreover, existing agentic search benchmarks (Table 1) focus on general domains and miss critical local-life challenges such as geographic constraints and multi-service integration. Therefore, there is an urgent need for a multi-hop benchmark tailored to local life scenarios, accurately reflecting the complexity of real-world search tasks and fully challenging existing agentic search methods.
@@
To bridge these gaps, we first introduce LocalSearchBench, a comprehensive benchmark comprising over 1,300,000 data entries across 6 primary service categories, spanning 9 major cities in China. These entries are transformed from raw data into high-quality augmented and anonymized database. Based on these data, we propose a high-quality instruction synthesis method to design 900 multi-hop QA tasks across two intelligence levels. These multi-hop QA tasks are constructed through a multi-agent data generation process, requiring agents to perform complex reasoning across multiple information sources and reasoning steps.
@@
Building upon LocalSearchBench, we further develop LocalPlayground, an agentic search framework to evaluate LLM agents for local life service. Given the absence of appropriate benchmarks and the critical need for comprehensive evaluation systems, our approach addresses these challenges through comprehensive dataset construction, systematic evaluation framework and targeted model evaluation. We build a comprehensive evaluation environment with LocalRAG based on merchant database and web search tools to simulate real-world local life service search scenarios where models need to retrieve and synthesize information from multiple sources, and conduct evaluations of multiple LRMs.
@@ 2. Related Work
Table 1. Comparison of existing user interaction benchmarks across key evaluation dimensions. “” indicates fully addressed, “ ” indicates partially addressed, and “” indicates not addressed.
Benchmark Reasoning Complexity Domain Specificity Task Complexity
Multi-hop Reasoning Tool Integration Context Dependency Domain Knowledge Geographic Constraints Multi-service Integration Real-world Scenarios Scale (samples)
HotpotQA (Yang et al., 2018) 113K
MuSiQue (Trivedi et al., 2022) 25K
BrowseComp (Wei et al., 2025) 1K
DeepResearch (Du et al., 2025) 500
DeepWideSearch (Lan et al., 2025a) 220
ResearchY (Rosset et al., 2024) 1.5K
ProxyQA (Tan et al., 2024) 1550
LocalSearchBench (ours) 1.35M + 900
@@ 2.1. Agentic Search
Gou et al. (2025) define agentic search as systems capable of integrating multiple auxiliary tools to autonomously perform interactive generation through search tools (e.g., web search APIs or external databases) for handling complex search tasks. Leveraging the powerful language understanding capabilities of LLMs, these search agents can proactively address ambiguous or underspecified queries through clarification or reasonable assumptions (Yang et al., 2025b). They decompose the initial search task into manageable sub-queries, perform dynamic reasoning and planning based on accumulated contextual information, adjust search strategies in real time, and synthesize information from multiple sources (Li et al., 2025c).
Recent advances in the reasoning abilities of LRMs further foster the development of LRM-based agentic search systems. The most representative products include OpenAI Deep Research (Jaech et al., 2024), Gemini Deep Research (Comanici et al., 2025), and Tongyi Deep Research (Team, 2025). In addition to accessing real-time information through search APIs, these systems leverage various auxiliary tools to enable deeper and broader research on highly complex problems. In the open-source community,  Li et al. (2025a) first proposed integrating the agentic search workflow into the o1-style reasoning process of LRMs to achieve autonomous knowledge supplementation.  Li et al. (2025b),  Jin et al. (2025), and  Song et al. (2025) further enhanced the autonomous search capabilities of LLMs by applying reinforcement learning, enabling the models to generate queries and invoke external retrieval systems during reasoning.
@@ 2.2. Benchmarking Agentic capabilities
Existing benchmarks for agentic search systems in general domain can be broadly categorized into closed-ended and open-ended QA tasks. Closed-ended QA benchmarks include multi-hop reasoning datasets that require synthesizing information from multiple sources (Yang et al., 2018; Ho et al., 2020; Press et al., 2022; Trivedi et al., 2022), challenging QA benchmarks designed for long-horizon questions with long-tail knowledge (Wei et al., 2025; Zhou et al., 2025; Xi et al., 2025; Pham et al., 2025), and fact-checking tasks that evaluate claim verification abilities (Wei et al., 2024; Wadden et al., 2020; Jiang et al., 2020; Wang et al., 2024). Open-ended QA benchmarks focus on deep information seeking tasks, targeting multi-perspective queries (Rosset et al., 2024) and expert-level research tasks (Du et al., 2025; Bosse et al., 2025; Tan et al., 2024), with some incorporating multi-modal capabilities (Yuan et al., 2025; Dong et al., 2025; Liu et al., 2025c).
@@ 3. Construction of LocalSearchBench
Figure 2. Workflow of Local Merchant Database Construction.
@@ 3.1. Local Merchant Database Construction
In this subsection, we present the construction of the local merchant database for LocalSearchBench, which serves as the foundation for multi-hop QA tasks. As shown in Figure 2, our database construction consists of three stages: Merchant Seed Data Collection, Data Augmentation and Synthesis, and Data Validation.
@@ 3.1.1. Merchant Seed Data Collection
To ensure LocalSearchBench reflects real-world local life service complexities, we built our local merchant database based on raw data from a leading local life service platform M. We sampled over 1,600,000 real merchant records spanning from January to December 2025, covering diverse service categories and landmarks across 9 major Chinese cities, involving over 9,900,000 retrievals from raw data, with the generation of these seed data requiring 1,418 hours of computation time.
@@
Based on platform M’s core business verticals, we identified six primary service categories reflecting real-world user demand: Dining (35%), Lifestyle (25%), Shopping (20%), Accommodation (10%), Healthcare (5%), and Tourism (5%), as shown in Figure 4(4(a)). We curated a keyword pool of 927 unique terms across these categories to reflect authentic user search behaviors and regional preferences.
@@
The collection spans 9 major Chinese cities with comprehensive coverage of benchmarks (Table 2). The location database includes 10,741 landmarks across diverse categories: commercial centers, transportation hubs, cultural landmarks, educational institutions, medical facilities, and residential areas. Figure 3 shows word clouds of landmark distributions for three of the nine cities.
Table 2. Geographic coverage across 9 major Chinese cities
City Landmarks Districts Retrievals
Shanghai 1556 16 1,442,412
Beijing 1451 16 1,345,077
Guangzhou 1026 11 951,102
Shenzhen 1039 9 963,153
Hangzhou 1129 13 1,046,583
Chengdu 1073 20 994,671
Chongqing 1008 38 934,416
Wuhan 1040 13 964,080
Suzhou 1149 9 1,065,123
Total 10,741 145 9,956,907
Merchant Seed Data Retrieval Query Generation
We construct scenario-specific queries by randomly combining 1–2 keywords from the 927-term pool to simulate natural search patterns. Applied across 145 administrative divisions, this multi-city keyword strategy generates diverse query variations and comprehensive coverage of local search intents and regional preferences, forming the foundation for LocalSearchBench’s merchant database synthesis.
@@ 3.1.2. Data Augmentation and Synthesis
The original merchant seed data contains only 12 basic fields with many missing or unstandardized attributes, which is inadequate for agentic search tasks that require dense, high-coverage merchant profiles. We employ a Data Augmentation Agent to expand these 12-field records into 29-field profiles using LLMs across six dimensions (see Figure 2 and Figure 14). The agent enriches each merchant’s basic, locational, operational, service, contact, and promotional information through 17 newly added fields, as illustrated in Figure 2.
@@
To comply with data protection regulations and enable safe public release, we employ a Privacy Rewriting Agent on the augmented data (see Figure 15). Real-world merchant data contains personally identifiable information and business-sensitive details that must be anonymized (Liu et al., 2023; Liu, 2024). The agent rewrites seven privacy-related fields to anonymize merchant identity, location, and contact details while keeping the remaining business information usable, as illustrated in Figure 2.
@@ 3.1.3. Data Validation
It evaluates augmented data quality against the original seed data along six weighted dimensions: Data Completeness (20%), Logical Consistency (30%), Geographic Accuracy (10%), Content Quality (20%), Format Standardization (10%), and Innovation Degree (10%), as summarized in Table 2(3).
Validation Thresholds: Thresholds are calibrated from an empirical study on 1,000 samples (overall score 0.86), benchmark dataset standards (minimum 0.80), and alignment with human ratings. A merchant record passes if the Overall Score $\geq 0.80$, all six dimensions $\geq 0.70$, key dimensions (Logical Consistency, Data Completeness, Content Quality) meet their target thresholds, and there are no critical failures in essential fields (name format, category alignment, address consistency).
@@
Privacy Rewriting Data Validation
It evaluates the quality of privacy rewriting against the augmented data along four dimensions: Anonymization Degree (50%), Logical Consistency (20%), Content Consistency (20%), and Innovation Degree (10%), as detailed in Table 2(3).
Validation Thresholds: Due to the higher privacy risk, this stage adopts stricter criteria informed by GDPR requirements and an empirical study on 1,000 samples (overall score 0.92, Anonymization Degree 0.95) (Liu et al., 2025b). A record passes if the Overall Score $\geq 0.85$, Anonymization Degree $\geq 0.90$, Content Consistency $\geq 0.90$, all dimensions $\geq 0.75$, and there are no privacy leaks in sensitive fields (name, address, email).
Table 4. Validation Thresholds Summary
Validation Stage Overall Score Critical Dimensions All Dimensions
Augmented Data $\geq$ 0.80
Logical Consistency $\geq$ 0.85
Data Completeness $\geq$ 0.80
Content Quality $\geq$ 0.80
$\geq$ 0.70
Privacy Rewriting $\geq$ 0.85
Anonymization Degree $\geq$ 0.90
Content Consistency $\geq$ 0.90
$\geq$ 0.75
@@ 3.1.4. Local Merchant Database Statistics
Overall, we constructed 1,354,185 representative samples from a pool of over 1,600,000 original seed records through data augmentation, privacy rewriting, and validation, which constitute our LocalSearchBench database.
The proportions of each category, the number of data for each city are presented in Figure 4(4(c))–4(4(k)). The distribution closely mirrors that of the raw data, ensuring our benchmark maintains the same categorical balance as real-world local life service platforms. The geographical distribution of merchant data for each city was derived through geospatial inference to generate latitude-longitude heatmaps, as shown in Figure 5. The distribution patterns align well with urban clustering phenomena and regional economic hierarchies, demonstrating the dataset’s realistic geographical representation.
(a) Standardized category distribution of real-world data
(b) City distribution
(c) Shanghai
(d) Beijing
(e) Guangzhou
(f) Shenzhen
(g) Hangzhou
(h) Chengdu
(i) Chongqing
(j) Wuhan
(k) Suzhou
Figure 4. City distribution and category distribution of 1,354,185 merchant data across 9 major cities in China.
@@
(d) Shenzhen
(e) Hangzhou
(f) Chengdu
(g) Chongqing
(h) Wuhan
(i) Suzhou
Figure 5. Geographical distribution heatmaps of 1,354,185 merchant data across 9 major cities in China. The color scale represents merchant density from 50 (blue) to 100 (red).